OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -5x + 15 -------- 5

OpenStudy (lalaly):

-x+3

OpenStudy (akshay_budhkar):

hi amistre

OpenStudy (amistre64):

hello

OpenStudy (akshay_budhkar):

truthfully even i am fed up with algebra now :)

OpenStudy (amistre64):

:) it can get rather repetitive

OpenStudy (akshay_budhkar):

it has got repetitive to me! need some good stuff now

OpenStudy (amistre64):

use 1/5 as the part to distribute thru the top

OpenStudy (amistre64):

or look at it as the distribution of division: \[\frac{-5x}{5}+\frac{15}{5}\] and reduce

OpenStudy (anonymous):

Like amistre said. Dividing by a number is the same as multiplying by its reciprocal, so rewrite that as (1/5) (-5x + 15) Then distribute and simplify the fractions.

OpenStudy (anonymous):

how about \[\frac{-5x+15}{5}=\frac{5(-x+3)}{5}=-x+3\]

OpenStudy (amistre64):

lol ... thats the undistributive property

OpenStudy (anonymous):

the old "factor an cancel" mantra

OpenStudy (anonymous):

Oooh. Gettin' fancy now, Satellite. Unfortunately you failed to follow the directions.

OpenStudy (amistre64):

bend then lift ...

OpenStudy (anonymous):

but they said "use the distributive property" right?

OpenStudy (anonymous):

Well I suppose you could say pulling out the 5 was distributive property.

OpenStudy (anonymous):

the distributive law says \[ab+ac=a(b+c)\] and that is what i used

OpenStudy (anonymous):

is -x + 3 also x - 3?

OpenStudy (lalaly):

3-x

OpenStudy (amistre64):

not unless you squeeze it into |..|

OpenStudy (anonymous):

Alright. I give you credit then.

OpenStudy (amistre64):

-5+3 = -2 5 - 3 = 2

OpenStudy (anonymous):

i only wrote it because i was pretty sure that is what the person who wrote the question had in mind

OpenStudy (anonymous):

guys your confusing me... please i dont get the answer

OpenStudy (akshay_budhkar):

OMG!! guys dont confuse him

OpenStudy (anonymous):

\[\frac{-5x + 15}{5} = (-5x + 15)\cdot 5^{-1} = ((-5\cdot5^{-1})x\ + 5^{-1}\cdot15 =-x + 3 \]

OpenStudy (amistre64):

the consensus is: -x+3 is the answer

OpenStudy (anonymous):

the distributive law says \[ab+ac=a(b+c)\] your numerator is \[-5x+15=-x(5)+3(5)=(-x+3)5\]

OpenStudy (anonymous):

another way to write the answer? x -3?

OpenStudy (anonymous):

now you can cancel the 5

OpenStudy (anonymous):

@andras no

OpenStudy (amistre64):

3-x is another way to write it if they dont like negatives to come first .. feng shuay stuff

OpenStudy (anonymous):

for example if x = 10 then \[10-3=7\] and \[3-10=-7\] one is the negative of the other

OpenStudy (akshay_budhkar):

i thought it was a very simple question :)

OpenStudy (anonymous):

\[\frac{-5x + 15}{5} = (-5x + 15)\cdot 5^{-1} =(-5\cdot5^{-1})x\ + 5^{-1}\cdot15 =-x + 3\]

OpenStudy (anonymous):

FEMG SHUI literally "new fad"

OpenStudy (anonymous):

Remember that division is really multiplying by the multiplicative inverse.

OpenStudy (amistre64):

..everybody wang chung tonite ;)

OpenStudy (akshay_budhkar):

feng shui! my mother tries to keep things according to it

OpenStudy (anonymous):

@amiste you are showing your age @akshay you used to be able to make money hiring yourself out as a feng shui consultant. i doubt you could these days